OpenStudy (anonymous):

19x^2-24x=0 x= Type an integer or simplified fraction. Use a comma to seperate answers as needed

OpenStudy (anonymous):

19x^2-24x=0 factor: x(19x-24) = 0 x = 0 or 19x-24 = 0 19x = 24 x = (24/19) x = 0 or 24/19

OpenStudy (anonymous):

so the answer is 24,19

OpenStudy (anonymous):

or 0,24/19

OpenStudy (anonymous):

24/19 yes 0 could also be an answer.

OpenStudy (anonymous):

theyre both answers actually